OpenStudy (andijo76):

can someone help me figure out this question i have no idea how to do this one the sum of two numbers is 56 the first number is 2/5 of the second number what are the numbers

OpenStudy (raden):

a+b = 56 ... (1) and a = 2/5 b ... (2) subtitute the 2nd equ to the 1st equ : a + b = 56 2/5 b + b = 56 times 5 to both sides, get 5 * 2/5 b + 5b = 56 * 5 2b + 5b = 280 7b = 280 both sides divided by 7 , get 7b/7 = 280/7 b = 40 to get the value of a, resubtitute b = 40 into the 2nd equation : a = 2/5 b so, a = 2/5 * 40 = 16 therefore, that number respectively are a = 16 and b = 40

OpenStudy (andijo76):

oh okay that makes sense to me thank you

OpenStudy (raden):

welcome